# Joint Cooperative and Non-Cooperative Localization in WSNs with Distributed Scaled Proximal ADMM Algorithms

### v1

@@ II-C Problem Reformulation
$\arg\min_{\mathbf{z},\mathbf{w}}$ $\sum_{i\in\mathcal{N}}\big[\underbrace{\frac{1}{2}\|\mathbf{H}_{i}\mathbf{z}_{i}\|^{2}-\mathbf{w}_{i}^{T}\mathbf{D}_{i}\mathbf{H}_{i}\mathbf{z}_{i}}_{G_{i}(\mathbf{z}_{i},\mathbf{w}_{i})}+\delta_{\mathcal{B}^{N_{i}+1}}\left(\mathbf{w}_{i}\right)\big]$ (24a)
@@
Figure 1: An illustrative example of a connected wireless sensor network in 2-D space with four sensors, one anchor, and one passive target.
To better understand these notations, an example of a wireless sensor network connectivity graph in two-dimensional space is illustrated in Fig. 1.
Example Fig. 1 shows a wireless sensor network connectivity graph consisting of three agents (with unknown positions), one anchor (with known position) and one target. Suppose we focus on sensor $2$ as a representative example, which attempts to localize both itself and the target. On one hand, since $\mathcal{N}_{2}=\{1,4\}$ and $\mathbf{p}_{2,1}^{-}$, $\mathbf{p}_{2,4}^{-}$ arevalues of $\mathbf{x}_{2}$; $\mathbf{q}_{2,1}^{-}$, $\mathbf{q}_{2,4}^{-}$ arevalues of $\mathbf{y}_{2}$, we have
$\mathbf{x}_{2}=\mathbf{p}_{2,1}^{-},\quad\mathbf{x}_{2}=\mathbf{p}_{2,4}^{-},$
$\mathbf{y}_{2}=\mathbf{q}_{2,1}^{-},\quad\mathbf{y}_{2}=\mathbf{q}_{2,4}^{-}.$
Moreover, the consensus condition requires that $\mathbf{q}_{2,1}^{+}$ and $\mathbf{q}_{2,4}^{+}$ equal $\mathbf{y}_{2}$, i.e.,
$\mathbf{y}_{2}=\mathbf{q}_{2,1}^{+},\quad\mathbf{y}_{2}=\mathbf{q}_{2,4}^{+}.$
On the other hand, Fig. 1 indicates that $\mathbf{z}_{2}$, $\mathbf{z}_{1}$ and $\mathbf{z}_{4}$ share some common elements. Specifically, in terms of sensor localization: $\mathbf{x}_{1}$ is an element of both $\mathbf{z}_{1}$ and $\mathbf{z}_{2}$; $\mathbf{z}_{2}$ and $\mathbf{z}_{4}$ also contain $\mathbf{x}_{4}$. In terms of target localization: $\mathbf{y}_{1}$ is an element of both $\mathbf{z}_{1}$ and $\mathbf{z}_{2}$; $\mathbf{z}_{2}$ and $\mathbf{z}_{4}$ also contain $\mathbf{y}_{4}$. To reflect this observation, the following constraints are introduced:
$\mathbf{p}_{2,1}^{+}=\mathbf{x}_{1}=\mathbf{p}_{1,2}^{+},\quad\mathbf{p}_{2,4}^{+}=\mathbf{x}_{4}=\mathbf{p}_{4,2}^{+},$
$\mathbf{q}_{2,1}^{+}=\mathbf{y}_{1}=\mathbf{q}_{1,2}^{+},\quad\mathbf{q}_{2,4}^{+}=\mathbf{y}_{4}=\mathbf{q}_{4,2}^{+}.$
The above equalities constitute a subset of the constraint set $\mathcal{Y}$ defined in Equation (22).
III The Proposed SP-ADMM-JCNL Algorithm
@@
$\mathcal{L}_{i}\left(\mathbf{z}_{i},\mathbf{w}_{i},\bm{\lambda}_{i}\right):=$ $G_{i}\left(\mathbf{z}_{i},\mathbf{w}_{i}\right)+\delta_{\mathcal{B}^{N_{i}+1}}\left(\mathbf{w}_{i}\right)+\left\langle\bm{\lambda}_{i},\mathbf{A}_{i}\mathbf{z}_{i}\right\rangle$
@@
and $\bm{\lambda}_{i}:=\text{vec}\left(\bm{\lambda}_{i,j}^{1},\bm{\lambda}_{i,j}^{2},\bm{\lambda}_{i,j}^{3},j\in N_{i}\right)\in\mathbb{R}^{3N_{i}n}$ represents the corresponding Lagrange multipliers and $c>0$ is a penalty parameter. The iterative update steps for the variables are obtained by applying the following SP-ADMM algorithm :
@@
To facilitate distributed implementation, we design the augmented Lagrangian function $\mathcal{L}$ such that it is separable with respect to each local variable $\mathbf{z}_{i}$, and remains convex when other variables are fixed. These structural properties enable efficient decentralized updates across networked nodes and support scalability in implementation.In each subproblem, scaled proximal terms are incorporated to improve numerical stability and support the convergence analysis. This treatment follows the methodology adopted in our earlier work and ensures that each subproblem remains strongly convex, despite the overall non-convexity of the joint formulation. Specifically, they ensure that each subproblem remains strongly convex and thus admits a unique solution:
@@
To see the first point, we reformulate the objective function of problem (26a). Combining (24a) and (III) and rearranging the quadratic term, we have
@@
$+\delta_{\mathcal{B}^{N_{i}+1}}\left(\mathbf{w}_{i}^{t}\right)+\frac{c}{2}\|\mathbf{z}_{i}\|_{\mathbf{B}_{i}^{T}\mathbf{B}_{i}}^{2},$ (27)
@@
Assuming that the selection matrix $\mathbf{B}_{i}^{T}\mathbf{B}_{i}$ is chosen such that $\mathbf{U}_{i}$ is a positive definite matrix, it follows from (III) that the objective function of subproblem (26a) is smooth and strongly convex, thereby ensuring the uniqueness of its solution. By completing the square, we obtain:
@@
$+\delta_{\mathcal{B}^{N_{i}+1}}(\mathbf{w}_{i}^{t+1})+\frac{c}{2}\|\mathbf{z}_{i}^{t}\|_{\mathbf{B}_{i}^{T}\mathbf{B}_{i}}^{2},$ (29)
@@
Substituting (III) into (26a) and omitting constant terms, we obtain:
@@
where $|\cdot|$ denotes element-wise absolute value.Based on the definitions of $\mathbf{A}_{i}$ and $\mathbf{H}_{i}$ in (II-C) and (II-C), we can easily obtain the specific form of matrix $c\mathbf{B}_{i}^{T}\mathbf{B}_{i}$. By substituting it into Equation (28), we ensure that $\mathbf{U}_{i}$ is a positive definite diagonal matrix of the form:
$\mathbf{U}_{i}=2\cdot\text{Diag}\Bigg(\Big[(c+1)N_{i}+1,c\cdot\mathbf{1}_{N_{i}},\mathbf{1}_{N_{i}},$
$2cN_{i}+1,c\cdot\mathbf{1}_{N_{i}},c\cdot\mathbf{1}_{N_{i}}\Big]\Bigg)$ $\otimes\mathbf{I}_{n}.$ (33)
@@
Furthermore, to provide an explicit expression for $\mathbf{z}_{i}^{t+1}$, similar to the definition of $\mathbf{z}_{i}$, we partition $\tilde{\mathbf{z}}_{i}:=\left[(\mathbf{x}_{i})^{T},(\tilde{\mathbf{p}}_{i}^{-})^{T},(\tilde{\mathbf{p}}_{i}^{+})^{T},(\mathbf{y}_{i})^{T},(\tilde{\mathbf{q}}_{i}^{-})^{T},(\tilde{\mathbf{q}}_{i}^{+})^{T}\right]^{T}$ into six components, denoted as $\tilde{\mathbf{z}}_{i}$, where:
$\tilde{\mathbf{p}}_{i}^{-}:=\text{vec}\left(\tilde{\mathbf{p}}_{i,j}^{-},j\in\mathcal{N}_{i}\right),\quad\tilde{\mathbf{p}}_{i}^{+}:=\text{vec}\left(\tilde{\mathbf{p}}_{i,j}^{+},j\in\mathcal{N}_{i}\right),$
$\tilde{\mathbf{q}}_{i}^{-}:=\text{vec}\left(\tilde{\mathbf{q}}_{i,j}^{-},j\in\mathcal{N}_{i}\right),\quad\tilde{\mathbf{q}}_{i}^{+}:=\text{vec}\left(\tilde{\mathbf{q}}_{i,j}^{+},j\in\mathcal{N}_{i}\right),$
@@
$\mathbf{x}_{i}^{t+1}$ $=\begin{cases}\tilde{\mathbf{x}}_{i}^{t+1},&\text{for }i\notin\mathcal{A}\\ \mathbf{a}_{i},&\text{for }i\in\mathcal{A}\quad,\end{cases}$ (34a)
$(\mathbf{p}_{i}^{-})^{t+1}$ $=\text{vec}\left(a_{1}(\tilde{\mathbf{p}}_{i,j}^{-})^{t+1}+a_{2}(\tilde{\mathbf{p}}_{j,i}^{+})^{t+1},j\in\mathcal{N}_{i}\right),$ (34b)
$(\mathbf{p}_{i}^{+})^{t+1}$ $=\text{vec}\left(a_{1}(\tilde{\mathbf{p}}_{i,j}^{+})^{t+1}+a_{2}(\tilde{\mathbf{p}}_{j,i}^{-})^{t+1},j\in\mathcal{N}_{i}\right),$ (34c)
$\mathbf{y}_{i}^{t+1}$ $=\tilde{\mathbf{y}}_{i}^{t+1},$ (34d)
$(\mathbf{q}_{i}^{-})^{t+1}$ $=\text{vec}\left(\frac{1}{2}\left((\tilde{\mathbf{q}}_{i,j}^{-})^{t+1}+(\tilde{\mathbf{q}}_{j,i}^{+})^{t+1}\right),j\in\mathcal{N}_{i}\right),$ (34e)
$(\mathbf{q}_{i}^{+})^{t+1}$ $=\text{vec}\left(\frac{1}{2}\left((\tilde{\mathbf{q}}_{i,j}^{+})^{t+1}+(\tilde{\mathbf{q}}_{j,i}^{-})^{t+1}\right),j\in\mathcal{N}_{i}\right),$ (34f)
where $a_{1}=\frac{c}{c+1},a_{2}=\frac{1}{c+1}.$
By examining equations (34a)–(34f), it can be observed that the update of $\mathbf{z}_{i}^{t+1}$ relies solely on the estimated values of its neighboring nodes $\tilde{\mathbf{z}}^{t+1},j\in\mathcal{N}_{i}$. Consequently, the variable updates can be performed in a distributed manner. The detailed proof is provided in Appendix A. For the optimization problem (26b), omitting the terms that are independent of $\mathbf{w}_{i}$, it can be simplified as
$\arg\min_{\mathbf{w}}\sum_{i\in\mathcal{N}}\bigg($ $-\langle\mathbf{H}_{i}^{T}\mathbf{D}_{i}\mathbf{w}_{i},\mathbf{z}_{i}^{t+1}\rangle+\delta_{\mathcal{B}^{N_{i}+1}}(\mathbf{w}_{i})$ (35)
@@
The above objective function is separable with respect to node $i$ and requires projection onto the unit ball. Therefore,
$\mathbf{w}_{i}^{t+1}=\text{proj}_{\mathcal{B}^{N_{i}+1}}\left(\mathbf{w}_{i}^{t}+\frac{1}{\rho}\mathbf{D}_{i}\mathbf{H}_{i}\mathbf{z}_{i}^{t+1}\right).$ (36)
@@
$\mathbf{w}_{i,k}^{t+1}=\frac{\mathbf{w}_{i,k}^{t}}{\max\{1,\|\mathbf{w}_{i,k}^{t}\|\}},\quad k\in\{1,...,N_{i},N_{i}+1\},$ (37)
where $\tilde{\mathbf{w}}_{i}^{t+1}=\text{vec}\left(\tilde{\mathbf{w}}_{i,k}^{t+1},k\in\{1,...,N_{i},N_{i}+1\}\right)=\mathbf{w}_{i}^{t}+\frac{1}{\rho}\mathbf{D}_{i}\mathbf{H}_{i}\mathbf{z}_{i}^{t+1}.$
@@
Input: Parameters $c$, $\rho$, initial values $\mathbf{z}_{i}^{0}$, $\mathbf{w}_{i}^{0}$, $\bm{\lambda}_{i}^{0}=\mathbf{0}$.
   Let $\mathbf{U}_{i}=2\cdot\text{Diag}\big(\big[(c+1)N_{i}+1,\ c\cdot\mathbf{1}_{N_{i}},\ \mathbf{1}_{N_{i}},$
    $2cN_{i}+1,\ c\cdot\mathbf{1}_{N_{i}},\ c\cdot\mathbf{1}_{N_{i}}\big]\big)\otimes\mathbf{I}_{n}$, for $i=1,\dots,N$.
@@
    Broadcast $\left(\tilde{\mathbf{z}}_{i,j}^{-}\right)^{t+1},\left(\tilde{\mathbf{z}}_{i,j}^{+}\right)^{t+1}$ to region $j\in\mathcal{N}_{i}$;
    Receive $\left(\tilde{\mathbf{z}}_{j,i}^{-}\right)^{t+1},\left(\tilde{\mathbf{z}}_{j,i}^{+}\right)^{t+1}$ from region $j\in\mathcal{N}_{i}$.

### v2

@@ 2.3 Problem Reformulation
$\arg\min_{\mathbf{z},\mathbf{w}}$ $\sum_{i\in\mathcal{N}}\big\{\underbrace{\frac{1}{2}\|\mathbf{H}_{i}\mathbf{z}_{i}\|^{2}-\mathbf{w}_{i}^{T}\mathbf{D}_{i}\mathbf{H}_{i}\mathbf{z}_{i}}_{G_{i}(\mathbf{z}_{i},\mathbf{w}_{i})}+\delta_{\mathcal{B}^{N_{i}+{\color[rgb]{0,0,0}M}}}\left(\mathbf{w}_{i}\right)\big\}$ (23a)
@@
3 The Proposed SP-ADMM-JCNL Algorithm
@@
$\mathcal{L}_{i}\left(\mathbf{z}_{i},\mathbf{w}_{i},\bm{\lambda}_{i}\right):=$ $G_{i}\left(\mathbf{z}_{i},\mathbf{w}_{i}\right)+\delta_{\mathcal{B}^{{\color[rgb]{0,0,0}N_{i}+M}}}\left(\mathbf{w}_{i}\right)+\left\langle\bm{\lambda}_{i},\mathbf{A}_{i}\mathbf{z}_{i}\right\rangle$
@@
In this formulation, $c>0$ is a penalty parameter and $\bm{\lambda}_{i}:=[\bm{\lambda}_{i}^{\text{sen}}\quad\bm{\lambda}_{i}^{\text{tar}}]\in\mathbb{R}^{(1+2M)N_{i}n}$represents the corresponding Lagrange multipliers, Specifically, $\bm{\lambda}_{i}^{\text{sen}}=\text{vec}\left(\lambda_{i,j}^{0},j\in N_{i}\right)\otimes\mathbf{I}^{n}$corresponds to the multipliers for the sensing constraints, while $\bm{\lambda}_{i}^{\text{tar}}=\text{vec}\left(\lambda_{i,j}^{k},j\in N_{i},k=1,2,...,2M\right)\otimes\mathbf{I}^{n}$ corresponds to those for the target constraints. The iterative update steps for the variables are obtained by applying the following SP-ADMM algorithm :
@@
To facilitate distributed implementation, we design the augmented Lagrangian function $\mathcal{L}$ such that it is separable with respect to each local variable $\mathbf{z}_{i}$, and remains convex when other variables are fixed. These structural properties enable efficient decentralized updates across networked sensors and support scalability in implementation.In each subproblem, scaled proximal terms are incorporated to improve numerical stability and support the convergence analysis. This treatment follows the methodology adopted in our earlier work and ensures that each subproblem remains strongly convex, despite the overall non-convexity of the joint formulation. Specifically, they ensure that each subproblem remains strongly convex and thus admits a unique solution:
@@
To see the first point, we reformulate the objective function of problem (25a). Combining (23a) and (3) and rearranging the quadratic term, we have
@@
$+\delta_{\mathcal{B}^{{\color[rgb]{0,0,0}N_{i}+M}}}\left(\mathbf{w}_{i}^{t}\right)+\frac{c}{2}\|\mathbf{z}_{i}\|_{\mathbf{B}_{i}^{T}\mathbf{B}_{i}}^{2},$ (26)
@@
Assuming that the selection matrix $\mathbf{B}_{i}^{T}\mathbf{B}_{i}$ is chosen such that $\mathbf{U}_{i}$ is a positive definite matrix, it follows from (3) that the objective function of subproblem (25a) is smooth and strongly convex, thereby ensuring the uniqueness of its solution. By completing the square, we obtain:
@@
$+\delta_{\mathcal{B}^{{\color[rgb]{0,0,0}N_{i}+M}}}(\mathbf{w}_{i}^{t+1})+\frac{c}{2}\|\mathbf{z}_{i}^{t}\|_{\mathbf{B}_{i}^{T}\mathbf{B}_{i}}^{2},$ (28)
@@
Substituting (3) into (25a) and omitting constant terms, we obtain:
@@
where $|\cdot|$ denotes element-wise absolute value. Based on the definitions of $\mathbf{A}_{i}$ and $\mathbf{H}_{i}$ in (19) and (17), we can easily obtain the specific form of matrix $c\mathbf{B}_{i}^{T}\mathbf{B}_{i}$. By substituting it into Equation (27), we ensure that $\mathbf{U}_{i}$ is a positive definite diagonal matrix of the form:
$\mathbf{U}_{i}=$
$2\cdot\text{Diag}\Bigg(\Big[{\color[rgb]{0,0,0}(c+1)N_{i}+M_{i}},c\cdot\mathbf{1}_{N_{i}},\mathbf{1}_{N_{i}},{\color[rgb]{0,0,0}2cN_{i}+\mathbb{I}_{\mathcal{M}_{i}}(1)},$
$c\cdot\mathbf{1}_{N_{i}},c\cdot\mathbf{1}_{N_{i}},...,{\color[rgb]{0,0,0}2cN_{i}+\mathbb{I}_{\mathcal{M}_{i}}(M)},c\cdot\mathbf{1}_{N_{i}},c\cdot\mathbf{1}_{N_{i}}\Big]\Bigg)\otimes\mathbf{I}_{n},$ (32)
where
$\forall k\in\mathcal{M},\quad\mathbb{I}_{\mathcal{M}_{i}}(k)=\begin{cases}1,&{\color[rgb]{0,0,0}k\in\mathcal{M}}_{i},\\ 0,&k\notin\mathcal{M}_{i}.\end{cases}$
@@
Furthermore, to provide an explicit expression for $\mathbf{z}_{i}^{t+1}$, similar to the definition of $\mathbf{z}_{i}$, we partition $\tilde{\mathbf{z}}_{i}:=[(\mathbf{x}_{i})^{T},(\tilde{\mathbf{p}}_{i}^{-})^{T},(\tilde{\mathbf{p}}_{i}^{+})^{T},(\mathbf{y}_{i,1})^{T},(\tilde{\mathbf{q}}_{i,1}^{-})^{T},(\tilde{\mathbf{q}}_{i,1}^{+})^{T},...,(\mathbf{y}_{i,M})^{T}$
$(\tilde{\mathbf{q}}_{i,M}^{-})^{T},(\tilde{\mathbf{q}}_{i,M}^{+})^{T}]^{T}$ into $1+M$ components
, where: $\tilde{\mathbf{p}}_{i}^{-}:=\text{vec}\left(\tilde{\mathbf{p}}_{i,j}^{-},j\in\mathcal{N}_{i}\right),\quad{\color[rgb]{0,0,0}\tilde{\mathbf{q}}_{i,k}^{-}:=\text{vec}\left(\tilde{\mathbf{q}}_{i,j,k}^{-},j\in\mathcal{N}_{i}\right)},$
$\tilde{\mathbf{p}}_{i}^{+}:=\text{vec}\left(\tilde{\mathbf{p}}_{i,j}^{+},j\in\mathcal{N}_{i}\right),\quad{\color[rgb]{0,0,0}\tilde{\mathbf{q}}_{i,k}^{+}:=\text{vec}\left(\tilde{\mathbf{q}}_{i,j,k}^{+},j\in\mathcal{N}_{i}\right).}$
@@
Sensor-related:
$\mathbf{x}_{i}^{t+1}$ $=\begin{cases}\tilde{\mathbf{x}}_{i}^{t+1},&i\notin\mathcal{A}\\ \mathbf{a}_{i},&i\in\mathcal{A},\end{cases}$ (33a)
$(\mathbf{p}_{i}^{-})^{t+1}$ $=\operatorname{vec}\bigl(c_{1}(\tilde{\mathbf{p}}_{i,j}^{-})^{t+1}+c_{2}(\tilde{\mathbf{p}}_{j,i}^{+})^{t+1},\ j\in\mathcal{N}_{i}\bigr),$ (33b)
$(\mathbf{p}_{i}^{+})^{t+1}$ $=\operatorname{vec}\bigl(c_{1}(\tilde{\mathbf{p}}_{i,j}^{+})^{t+1}+c_{2}(\tilde{\mathbf{p}}_{j,i}^{-})^{t+1},\ j\in\mathcal{N}_{i}\bigr),$ (33c)
where $c_{1}=\frac{c}{c+1},c_{2}=\frac{1}{c+1}.$
Target-related: $\forall{\color[rgb]{0,0,0}k}\in\mathcal{M},$
$\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}^{t+1}$ $=\tilde{\mathbf{y}}_{i,{\color[rgb]{0,0,0}k}}^{t+1},$ (33d)
$(\mathbf{q}_{i,{\color[rgb]{0,0,0}k}}^{-})^{t+1}$ $=\operatorname{vec}\left(\frac{1}{2}\bigl((\tilde{\mathbf{q}}_{i,j,{\color[rgb]{0,0,0}k}}^{-})^{t+1}+(\tilde{\mathbf{q}}_{j,i,{\color[rgb]{0,0,0}k}}^{+})^{t+1}\bigr),\ j\in\mathcal{N}_{i}\right),$ (33e)
$(\mathbf{q}_{i,{\color[rgb]{0,0,0}k}}^{+})^{t+1}$ $=\operatorname{vec}\left(\frac{1}{2}\bigl((\tilde{\mathbf{q}}_{i,j,{\color[rgb]{0,0,0}k}}^{+})^{t+1}+(\tilde{\mathbf{q}}_{j,i,{\color[rgb]{0,0,0}k}}^{-})^{t+1}\bigr),\ j\in\mathcal{N}_{i}\right).$ (33f)
By examining equations (33a)–(33f), it can be observed that the update of $\mathbf{z}_{i}^{t+1}$ relies solely on the estimated values of its neighboring sensors $\tilde{\mathbf{z}}^{t+1},j\in\mathcal{N}_{i}$. Consequently, the variable updates can be performed in a distributed manner. The detailed proof is provided in See Appendix A.1. For the optimization problem (25b), omitting the terms that are independent of $\mathbf{w}_{i}$, it can be simplified as
$\arg\min_{\mathbf{w}}\sum_{i\in\mathcal{N}}\bigg($ $-\langle\mathbf{H}_{i}^{T}\mathbf{D}_{i}\mathbf{w}_{i},\mathbf{z}_{i}^{t+1}\rangle+\delta_{\mathcal{B}^{{\color[rgb]{0,0,0}N_{i}+M}}}(\mathbf{w}_{i})$ (34)
@@
The above objective function is separable with respect to sensor $i$ and requires projection onto the unit ball. Therefore,
$\mathbf{w}_{i}^{t+1}=\text{proj}_{\mathcal{B}^{{\color[rgb]{0,0,0}N_{i}+M}}}\left(\mathbf{w}_{i}^{t}+\frac{1}{\rho}\mathbf{D}_{i}\mathbf{H}_{i}\mathbf{z}_{i}^{t+1}\right).$ (35)
@@
$\mathbf{w}_{i,l}^{t+1}=\frac{\mathbf{w}_{i,l}^{t}}{\max\{1,\|\mathbf{w}_{i,l}^{t}\|\}},\quad l\in\{1,...,N_{i}+M\},$ (36)
where $\tilde{\mathbf{w}}_{i}^{t+1}=\text{vec}\left(\tilde{\mathbf{w}}_{i,l}^{t+1},l\in\{1,...,N_{i}+M\}\right)=\mathbf{w}_{i}^{t}+\frac{1}{\rho}\mathbf{D}_{i}\mathbf{H}_{i}\mathbf{z}_{i}^{t+1}.$
@@
Input: Parameters $c$, $\rho$, initial values $\mathbf{z}_{i}^{0}$, $\mathbf{w}_{i}^{0}$, $\bm{\lambda}_{i}^{0}=\mathbf{0},$
    $\forall i=1,\dots,N$. Let $\mathbf{U}_{i}$ be as in (3).
@@
    Broadcast $(\tilde{\mathbf{p}}_{i,j}^{-})^{t+1},(\tilde{\mathbf{q}}_{i,j,{\color[rgb]{0,0,0}k}}^{-})^{t+1},k\in\mathcal{M}$ to $j\in\mathcal{N}_{i}$;
    Receive $(\tilde{\mathbf{p}}_{j,i}^{-})^{t+1},(\tilde{\mathbf{q}}_{j,i,k}^{+})^{t+1},k\in\mathcal{M}$ from $j\in\mathcal{N}_{i}$.